Option Informatique en Spé MP^{∗} — Année scolaire 2005/2006

1 ABERKANE Moad
2 ALVAREZ Sebastien
3 ASSEL Benjamin*
4 AUBERT Nathanaël
5 BELKABIR Yassine*
6 BENASSI Romain*
7 BENUREAU Fabien*
8 BOUHAFID Yassine
9 CHOU Remi
10 GABARROU Sylvain*
11 GILLET Guillaume*
12 GOMES Sebastien
13 GUIRAUD Venceslas
14 KEITA Doudou
15 KELLOU Nabil
16 LANCIEN Nicolas
17 PIN Furcy
18 PROVOST Nicolas
19 SICSIC Adrien
20 SINGER Amandine